# Single vibronic level fluorescence spectra from Hagedorn wavepacket dynamics

### v1

@@ III.1 Two-dimensional harmonic potentials
In higher dimensions, the quantum split-operator approach quickly becomes unfeasible due to the curse of dimensionality, particularly when treating excited vibrational wavefunctions for which denser grids are generally needed. In contrast, Hagedorn wavepackets circumvent the need for a grid and can be efficiently propagated within a harmonic approximation using the same trajectory as the simple thawed Gaussian. With the algebraic expressions developed in Ref. 49, we can now use Hagedorn wavepackets in much higher dimensions.
Figure 3 shows SVL emission spectra in a 100-dimensional system that incorporates displacement, distortion, and Duschinsky rotation effects. Except for the potential energy surface and initial wavepackets, the computation parameters (time step, total time of propagation, frequency of evaluation of the autocorrelation function, broadening of the spectra) are identical to those for the two-dimensional systems. We choose two modes (namely modes 30 and 40) to be vibrationally excited in the initial wavepackets, computing emission spectra from initial vibrational levels $30^{a}40^{b}$, with the same set of $(a,b)$ as in the two-dimensional cases.
@@
The spectra become significantly more complex in higher-dimensional systems. We observe broadened tails in the lower frequency region, which are likely attributable to the effects of Duschinsky rotation here (compare to Fig. 4). This also demonstrates the advantage of the time-dependent approach, where Duschinsky rotation can be intrinsically taken into account, in contrast to the time-independent method, which faces challenges related to the pre-selection of likely transitions and the computation of non-separable Franck-Condon integrals in high-dimensional systems.
Figure 4: SVL emission spectra from different vibrational levels (indicated in each panel) in a 100-dimensional two-state displaced and distorted harmonic system (without Duschinsky rotation).

### v2

@@ III.1 Two-dimensional harmonic potentials
In higher dimensions, the quantum split-operator approach quickly becomes unfeasible due to the curse of dimensionality, particularly when treating excited vibrational wavefunctions for which denser grids are generally needed. In contrast, Hagedorn wavepackets circumvent the need for a grid and can be efficiently propagated within a harmonic approximation using the same trajectory as the simple thawed Gaussian. With the algebraic expressions developed in Ref. 67, we can now use Hagedorn wavepackets in much higher dimensions.
Figure 4 shows SVL emission spectra in a 100-dimensional system that incorporates displacement, distortion, and Duschinsky rotation effects (all parameters are specified in the supplementary material). Except for the potential energy surface and initial wavepackets, the computation parameters (time step, total time of propagation, frequency of evaluation of the autocorrelation function, broadening of the spectra) are identical to those for the two-dimensional systems. We choose two modes (namely modes 30 and 40) to be vibrationally excited in the initial wavepackets, computing emission spectra from initial vibrational levels $30^{a}40^{b}$, with the same set of $(a,b)$ as in the two-dimensional cases.
@@
The spectra become significantly more complex in higher-dimensional systems. In this particular example, similarity can be observed between the $30^{0}40^{0}$ and $30^{1}40^{0}$ spectra as well as between the $30^{1}40^{1}$ and $30^{2}40^{0}$ spectra. However, significant changes in the intensity pattern can be seen when comparing spectra with different excitations in mode 40 (compare $30^{0}40^{0}$ and $30^{0}40^{3}$, or $30^{1}40^{0}$ and $30^{1}40^{1}$), indicating that the excitation in mode 40 likely has a more pronounced impact on the overall spectrum compared to the excitation in mode 30. We also observe broadened tails in the lower frequency region, which are likely attributable to the effects of Duschinsky rotation here (compare to Fig. 5 where the Duschinsky couplings are neglected). As a time-dependent approach, Hagedorn wavepacket dynamics can intrinsically take into account Duschinsky rotation effects, in contrast to the time-independent method, which faces challenges related to the pre-selection of likely transitions and the computation of non-separable Franck-Condon integrals in high-dimensional systems.
Figure 5: SVL emission spectra from different vibrational levels (indicated in each panel) in a 100-dimensional two-state displaced and distorted harmonic system (same as in Fig. 4 but without Duschinsky rotation).
@@ Acknowledgements.
Author declarations
Conflict of Interest
The authors have no conflicts to disclose.
Data availability
The data that supports the findings of this study are available within the article and its supplementary material.